For the equation y=1/2x+3 what numbers do I put in the y and x axis
Answer by schmidtmomto6(24)

In y=1/2x+3, 1/2 is the slope of the line, and 3 is the Y intercept. To find where the line intercepts the x axis, (which is where the Y value is 0) you just plug in a 0 for the Y and solve the equation, like this: 0=1/2x+3; then solve for x. So, -3=X/2, multiply each side times 2, then you have -6=X.So the line intercepts the X axis at point (-6,0) and intercepts the y axis at point (0,3)